Aggregator
Hall B current Oscillations
Log Number: 4370122Logbook(s): Entry Makers: shookBody:
I noticed that Hall B's current is fluctuating a lot this evening after the Injector resetup. While the HALO rates haven't changed muched, improved a little even, the current is jumping around much more than it was last night. The gun is also pushing much more current (~13uA->~20uA) to make their 325nA and the attenuator is much higher than it was previously (~400->~470). It also appears to oscillate a bit more as well. Not sure if this is a big issue, but wanted to make note of it.
MBC9R02V mismatch
Emittance Measurement: IHA1C07A X-plane - ABORTED
HYST AREA 6-4-1 Cycles Magnets for eDT
eDT V2-36 Design Setpoints Download
HarpFitter: IHA1C12
eDT V2-36 Design Setpoints Download
HarpFitter: IHA1C05
HarpFitter: IHA1C05
Operator SIGN IN
FSD Masking Tool: Successful System Masking for: BeamLossMonitor ***Some node's masks failed to write
FSD Masking Tool: Successful System UN-Masking for: BeamLossMonitor ***Some node's masks failed to write
Log Number: 4370103Logbook(s): Tags: Body: System UN-Masking for: BeamLossMonitor
User: shook
Areas: All
BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T
User: shook
Areas: All
BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T
Fopt 25may060243 results
Log Number: 4370102Logbook(s): Body: Fopt results from /usr/ntfs/public/optics/diforb/25may060243
Comment:
HOSTNAME: opsl01
CED WORKSPACE: OPS CED LOGIN KEY: pro
MYA DEPLOYMENT: ops
CONFIG PATH: /cs/prohome/apps/q/QTSnyder_QT5/1-1/fileio/config/cs.conf
DIFORB PATH: /usr/ntfs/public/optics/diforb
Comment:
HOSTNAME: opsl01
CED WORKSPACE: OPS CED LOGIN KEY: pro
MYA DEPLOYMENT: ops
CONFIG PATH: /cs/prohome/apps/q/QTSnyder_QT5/1-1/fileio/config/cs.conf
DIFORB PATH: /usr/ntfs/public/optics/diforb
Fopt - Hall A check
Log Number: 4370101Logbook(s): Tags: Body: Fopt 7.9 completed successfully. The runtime settings include a sample count of 15, a minimum current of 4, and an operating regime of Standard with Excitation in 1S Region. The corrector and RF kicks applied were 1.50003 and 1.50003 volts respectively. The max. amplitude setting was 1 millimeters. The files for this run were saved in the directory /cs/opshome/hlapps/optics/diforb/25may060243. The X and Y stability of the machine during the run was 0.0544544 and 0.0436055 (mm) respectively.
FSD Masking Tool: Successful System Masking for: BeamLossMonitor ***Some node's masks failed to write
Log Number: 4370100Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: shook
Areas: All
BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09<
User: shook
Areas: All
BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09<
Autolog save
iocsoftw3spce restored
Log Number: 4370097Logbook(s): Tags: Body: restorelib 2.6
Target restore date was 2025-05-06 02:21:38
The output from the myRestore execution was as follows:
iocsoftw3spce
systemEnum@2025-05-06 02:21
2(2) restored
Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -i iocsoftw3spce (success)
PVs modified: 0
PVs modified: 0
PVs modified: 0
PVs modified: 0
PVs modified: 0
PVs modified: 0
Procedure return status: Success
HYST AREA 6-4-1 Cycles Magnets for eDT
eDT V2-36 Design Setpoints Download
Log Number: 4370094Logbook(s): Tags: Body: eDT V2-36 Design Setpoints Download
hostname: opsl03
CED workspace: OPS
CED login_key: pro
CHANNELNEW VALUEOLD VALUEMQK1C05.BDL-71582.6474-71582.6484MQK1C07.BDL-53419.8411-40419.0000MQK1C08.BDL53419.831438419.0000MQR1C19.BDL-4380.7755-2880.7744MQR1C20.BDL6868.70325368.7031
hostname: opsl03
CED workspace: OPS
CED login_key: pro
CHANNELNEW VALUEOLD VALUEMQK1C05.BDL-71582.6474-71582.6484MQK1C07.BDL-53419.8411-40419.0000MQK1C08.BDL53419.831438419.0000MQR1C19.BDL-4380.7755-2880.7744MQR1C20.BDL6868.70325368.7031